The question is write for the line with the points (4,5) (6,9). i'm not sure how to figure it out, although I know it has to be in the form of y=mx+b
Answer by rfer(16322)

first find slope
m=(9-5)/(6-4)
m=4/2
m=2
---------------
then fin y int
9=2*6+b
9=12+b
-3=b
----------------
y=2x-3